Beale Manufacturing Company has a beta of 1.4, and Foley Industries has a beta of 0.5. The required return on an index fund that holds the entire stock market is 8%. The risk-free rate of interest is 3%. By how much does Beale's required return exceed Foley's required return? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Required return=risk free rate+beta*(market rate-risk free rate)

Required return for:

Beale Manufacturing Company=3+(8-3)*1.4=10%

Foley Industries=3+(8-3)*0.5=5.5%

Hence difference=10-5.5

=4.5%